# Trace field degrees in the Torelli group

### v1

@@ 3. An irreducibility criterion
The goal of this section is to present an algebraic criterion that allows us to deduce that certain characteristic polynomials of matrices of the form $XX^{\top}$ are irreducible.
@@
Let $M$ be a square integer matrix, and let $N$ be the principal submatrix of $M$ obtained by deleting the first row and the first column. If $M$ and $N$ have no common eigenvalue, and if $M$ has a simple eigenvalue $\rho$, then the characteristic polynomial of $\widetilde{M}=M+ay^{p}E_{11}$ is an irreducible element of $\mathbb{Z}[t,y]$, for all $p\geq 1$ and for all $0\neq a\in\mathbb{Z}$.
@@
Our goal is to use Eisenstein’s criterion on $\chi_{\widetilde{M}}\in\mathbb{Z}[t,y]\cong\left(\mathbb{Z}[t]\right)[y]$, viewing it as a polynomial in the variable $y$ and coefficients in $\mathbb{Z}[t]$. We calculate
$\chi_{\widetilde{M}}(t,y)=\det(t\cdot\mathrm{Id}-{\widetilde{M}})=-y^{p}a\chi_{N}(t)+\chi_{M}(t)$
and notice that $a\chi_{N}$ and $\chi_{M}$ are relatively prime in $\mathbb{Z}[t]$. Indeed, $\chi_{M}$ has leading coefficient $+1$ and no root in common with $\chi_{N}$ by our assumption that $M$ and $N$ have no eigenvalue in common. In particular, they have no common factor, which shows that $\chi_{\widetilde{M}}\in\left(\mathbb{Z}[t]\right)[y]$ is primitive. In order to apply Eisenstein’s criterion, let $\mu_{\rho}\in\mathbb{Z}[t]$ be the minimal polynomial of the simple eigenvalue $\rho$ of $M$. By assumption, $\mu_{\rho}$ divides $\chi_{M}$ exactly once, but it does not divide $\chi_{N}$ since $\chi_{M}$ and $\chi_{N}$ have no common root. In particular, Eisenstein’s criterion applies to show that the polynomial $\chi_{\widetilde{M}}\in\left(\mathbb{Z}[t]\right)[y]\cong\mathbb{Z}[t,y]$ is irreducible. ∎
@@
Oscillatory matrices satisfy a stronger version of Perron-Frobenius theory, namely all the eigenvalues are positive real, simple, and they strictly interlace when taking a principal submatrix [And87]. Hence, 6 applies very cleanly to this class of matrices.
We use 6 on the following two cases (9 and 10).
@@
be square integer matrices with $a_{1}\geq 1$. If $M$ is nonnegative and irreducible, and if $\chi_{N}\in\mathbb{Z}[t]$ is irreducible, then the characteristic polynomial of $\widetilde{M}=M+ay^{2}E_{11}$ is irreducible in $\mathbb{Z}[t,y]$ for all $0\neq a\in\mathbb{Z}$.
@@
In order to use 6, we need to show that $M$ has a simple eigenvalue and that $M$ and $N$ share no eigenvalue. The former holds since $M$ is nonnegative and irreducible, and in particular has a Perron-Frobenius eigenvalue which is simple. For the latter, we compute
@@
where $q(t)\in\mathbb{Z}[t]$ is of degree at most $n-1$. We claim that it is not the zero polynomial either. Indeed, we directly verify
@@
Now if there existed a common root $\lambda\in\mathbb{C}$ of $\chi_{M}$ and $\chi_{N}$, then $\lambda$ would also be a root of $q(t)$. But since $\chi_{N}$ is irreducible of degree $n$ and $q(t)$ is a nonzero polynomial of degree at most $n-1$, this is impossible. ∎
@@
be square integer matrices of dimension $n$ and $m$, respectively, with $a_{1},b_{1}\geq 1$. Furthermore, let $\alpha,\beta\neq 0$ such that
@@
is a matrix with integer coefficients. If $M$ is nonnegative and irreducible, and if $\chi_{A},\chi_{B}\in\mathbb{Z}[t]$ are irreducible and distinct, then the characteristic polynomial of $\widetilde{M}=M+ay^{2}E_{11}$ is irreducible in $\mathbb{Z}[t,y]$ for all $0\neq a\in\mathbb{Z}$.
@@
The proof is similar to the proof of 9: the only thing to verify is that no eigenvalue of $A$ or of $B$ is also an eigenvalue of $M$. Again, we compute
@@
where $q_{1}(t)\in\mathbb{Z}[t]$ is of degree at most $n-1$ and $q_{2}(t)\in\mathbb{Z}[t]$ is of degree at most $m-1$. This is seen by developing the first column of the matrix $tI-M$. The first coefficient is responsible for the summand $t\chi_{A}(t)\chi_{B}(t)$, the next $n$ coefficients are responsible for the summand ±$\pm q_{1}(t)\chi_{B}(t)$ and the final $m$ coefficients are responsible for the summand ±$\pm q_{2}(t)\chi_{A}(t)$. We claim that neither among $q_{1}(t)$ and $q_{2}(t)$ is the zero polynomial. Indeed, by developing the first column of the matrix $t\mathrm{I}-M$, and evaluating at $t=0$, we get
@@
which is not zero since $\chi_{A}$ is irreducible. Similarly, $q_{2}(0)\neq 0$. Now if there existed a common root $\lambda\in\mathbb{C}$ of $\chi_{M}$ and $\chi_{A}$, then $\lambda$ would also be a root of either $\chi_{B}$ or $q_{1}$. Since $\chi_{A}$ and $\chi_{B}$ are irreducible and distinct, we must have $q_{1}(\lambda)=0$. But since $\chi_{A}$ is irreducible of degree $n$, and $q_{1}(t)$ is a nonzero polynomial of degree at most $n-1$, this is impossible. Similarly, no root of $\chi_{B}$ can be a root of $\chi_{M}$, which concludes the proof. ∎
@@
One could formulate 10 with $k\geq 2$ blocks $A_{1},\dots,A_{k}$ of respective sizes $n_{1},\dots,n_{k}$, instead of $k=2$. In this case, all the $k$ characteristic polynomials $\chi_{A_{i}}$ should be irreducible and pairwise distinct. The argument is identical by considering
$\chi_{M}(t)=t\prod_{i=1}^{k}\chi_{A_{i}}+\sum_{i=1}^{k}\pm q_{i}(t)\prod_{j\neq i}\chi_{A_{j}},$
where $q_{i}(t)\in\mathbb{Z}[t]$ is of degree at most $n_{i}-1$ and nonzero.
@@ 4. Warm-up: proof of Thurston’s claim
As a first illustration of our method, it is the goal of this section to provide a pair of filling multicurves $\alpha$ and $\beta$ on $S_{g}$ with multicurve intersection degree $33g-$ 3. By 4, this validates Thurston’s claim that the product of two multitwists can realise the maximal possible algebraic degrees of stretch factors: $66g-$ 6.
@@
We start by realising, on the surface of genus $g\geq 1$ with $2$ boundary components, a pair of filling multicurves $\alpha$ and $\beta$ such that $\chi_{XX^{\top}}\in\mathbb{Z}[t]$ is irreducible and of degree $13g-$ 1. We proceed by induction on $g$.
@@
We consider the two multicurves $\alpha$ and $\beta$ shown in Figure 1, where one of the components of $\beta$ has $y-1$ parallel copies. Here, we think of $y$ as a variable that we specify later on.
$y-1$ copies
@@
Figure 1. Two multicurves $\alpha$ (in red) and $\beta$ (in blue) on the surface of genus one with two boundary components. The multicurve $\beta$ contains $y-1$ parallel copies of one of its components.
We directly calculate
@@
Observe that $X$ is a matrix of size $2\times y$ (the multicurve $\beta$ has $y$ components). We have $\chi_{XX^{\top}}(t)=t^{2}-(4+y)t+4(y-1)$ with discriminant $y^{2}-8y+32$, which is not a square if $y\geq 12$. Indeed, in this case we have
@@
For the inductive step, assume we have constructed on the surface of genus $g\geq 1$ with $2$ boundary components a pair of multicurves $\alpha^{\prime},\beta^{\prime}$ such that the characteristic polynomial $\chi^{\prime}=\chi_{XX^{\top}}\in\mathbb{Z}[t]$ is irreducible and of degree $13g-$ 1. Furthermore, assume that $\alpha^{\prime}_{1}$ is a simple closed curve that encircles all the handles of the surface, as illustrated in Figure 2. Take a surface of genus $1$ and two boundary components, as in the case of genus $g=1$, see Figure 1, and denote its multicurves by $\alpha^{\prime\prime}$ and $\beta^{\prime\prime}$. Now glue its right boundary component to the left boundary component of the surface of genus $g$, and add two new curves $\alpha_{0}$ and $\beta_{0}$ to the multicurves. The curve $\alpha_{0}$ encircles all the handles of the newly formed surface, and the curve $\beta_{0}$ twice intersects $\alpha_{0}$ but no other multicurve component. Again, see Figure 2 for an illustration.
@@
Figure 2. Two surfaces of genus $g$ and $1$, respectively, and two boundary components, glued together along one of their boundary components. The curves $\alpha^{\prime}_{1}$ and $\alpha^{\prime\prime}_{1}$ are shown, each encircling all the handles of their respective surface. The new curve $\alpha_{0}$ encircles all the handles of the newly formed surface, and the new curve $\beta_{0}$ runs along the glued boundary component.
@@
$\alpha$ $=\alpha_{0}\cup\alpha^{\prime}\cup\alpha^{\prime\prime}$
$\beta$ $=\beta_{0}\cup\beta^{\prime}\cup\beta^{\prime\prime}$
@@
Let us choose $b$ such that $\chi_{B}$ is irreducible and distinct from $\chi_{A}$. We may assume inductively that $a_{1}=4a$. In the multicurve $\beta$, we take $y^{2}-a-1\geq 1$ parallel copies of $\beta_{0}$, for $y>0$ large enough. The matrix $XX^{\top}$ for the multicurves $\alpha$ and $\beta$ takes the form
$XX^{\top}=\left(\begin{array}[]{c|ccc|cc}4y^{2}&a_{1}&\dots&a_{n}&4&2\\ \hline\cr a_{1}&&&&&\\ \vdots&&A&&&\\ a_{n}&&&&&\\ \hline\cr 4&&&&4&2\\ 2&&&&2&b\par\end{array}\right).$
[…]
$3g-1+3=3(g+1)-1,$
which is exactly what we wanted to show. Finally, to justify our inductive assumption on the top-left coefficient of the matrix $A$, note that the top-left coefficient of the matrix $XX^{\top}$ is again a multiple of 4.
@@
Take any example of a pair of multicurves $\alpha^{\prime}$ and $\beta^{\prime}$ we constructed on the surface of genus $g-1\geq 1$ with two boundary components in Section 4. Let
@@
be the matrix $XX^{\top}$ for the multicurves $\alpha^{\prime}$ and $\beta^{\prime}$, where $a_{1}=4a$. We identify the two boundary components of the surface to increase the genus by one. Let $\alpha_{0}$ be a longitude of the created handle, and let $\beta_{0}$ run along the glued boundary. Define the two new multicurves
@@
where we take $y^{2}-a$ copies of $\beta_{0}$. Then the matrix $XX^{\top}$ for the multicurves $\alpha$ and $\beta$ takes the form
@@
and $\chi_{XX^{\top}}\in\mathbb{Z}[t,y]$ is irreducible by 9. By Hilbert’s irreducibility theorem, there exist infinitely many specifications of $y$ (and in particular infinitely many specifications of $y$ such that $y^{2}-a>0$) with $\chi_{XX^{\top}}\in\mathbb{Z}[t]$ irreducible. This polynomial is of degree $33(g-1)-1+1=3g-$ 3 .

### v2

@@ 3. An irreducibility criterion
The goal of this section is to present an algebraic criterion that allows us to deduce that certain characteristic polynomials of matrices of the form $XX^{\top}$ are irreducible. We write $E_{11}$ for the matrix with the top-left coefficient equal to 11 and all other coefficients zero. Furthermore, for a square matrix $A$ with coefficients in a ring $R$, we denote its characteristic polynomial by $\chi_{A}=\det(t\cdot\mathrm{Id}-A)\in R[t]$.
@@
Let $M$ be a square integer matrix, and let $N$ be the principal submatrix of $M$ obtained by deleting the first row and the first column. If $M$ and $N$ have no common eigenvalue, and if $M$ has a simple eigenvalue $\rho$, then the characteristic polynomial of $\widetilde{M}=M+{c}y^{p}E_{11}$ is an irreducible element of $\mathbb{Z}[t,y]$, for all $0\neq p\in\mathbb{N}$ and $0\neq{c}\in\mathbb{Z}$.
@@
Our goal is to use Eisenstein’s criterion on $\chi_{\widetilde{M}}\in\mathbb{Z}[t,y]\cong\left(\mathbb{Z}[t]\right)[y]$, viewing it as a polynomial in the variable $y$ and coefficients in $\mathbb{Z}[t]$. By multilinearity of the determinant, we calculate
$\chi_{\widetilde{M}}(t,y)=\det(t\cdot\mathrm{Id}-{\widetilde{M}})=-y^{p}{c}\chi_{N}(t)+\chi_{M}(t)$
and notice that ${c}\chi_{N}$ and $\chi_{M}$ are relatively prime in $\mathbb{Z}[t]$. Indeed, $\chi_{M}$ has leading coefficient +1+1 and no root in common with $\chi_{N}$ by our assumption that $M$ and $N$ have no eigenvalue in common. In particular, they have no common factor, which shows that $\chi_{\widetilde{M}}\in\left(\mathbb{Z}[t]\right)[y]$ is primitive. In order to apply Eisenstein’s criterion, let $\mu_{\rho}\in\mathbb{Z}[t]$ be the minimal polynomial of the simple eigenvalue $\rho$ of $M$. By assumption, $\mu_{\rho}$ divides $\chi_{M}$ exactly once, but it does not divide $\chi_{N}$ since $\chi_{M}$ and $\chi_{N}$ have no common root. In particular, Eisenstein’s criterion applies to show that the polynomial $\chi_{\widetilde{M}}\in\left(\mathbb{Z}[t]\right)[y]\cong\mathbb{Z}[t,y]$ is irreducible. ∎
@@
Oscillatory matrices satisfy a stronger version of Perron–Frobenius theory, namely all the eigenvalues are positive real, simple, and they strictly interlace when taking a principal submatrix [And87]. Hence, 7 applies very cleanly to this class of matrices.
We use 7 on the following two cases (10 and 11).
@@
be square integer matrices with $a_{1}\geq 1$ and $0\neq\alpha\in\mathbb{Q}$. If $M$ is nonnegative and irreducible, and if $\chi_{N}\in\mathbb{Z}[t]$ is irreducible, then the characteristic polynomial of $\widetilde{M}=M+{c}y^{2}E_{11}$ is irreducible in $\mathbb{Z}[t,y]$ for all $0\neq{c}\in\mathbb{Z}$.
@@
In order to use 7, we need to show that $M$ has a simple eigenvalue and that $M$ and $N$ share no eigenvalue. The former holds since $M$ is nonnegative and irreducible, and in particular has a Perron–Frobenius eigenvalue which is simple, see for example [Gan59]. For the latter, we compute
@@
where $q(t)\in\mathbb{Z}[t]$ is of degree at most n1n-1. We claim that it is not the zero polynomial either. Indeed, we directly verify
@@
Now if there existed a common root $\lambda\in\mathbb{C}$ of $\chi_{M}$ and $\chi_{N}$, then $\lambda$ would also be a root of $q(t)$, as the equation $\chi_{M}(t)=t\chi_{N}(t)+q(t)$ implies $q(\lambda)=\chi_{M}(\lambda)-\lambda\chi_{N}(\lambda)=0$. Since $\chi_{N}$ is irreducible, it is the minimal polynomial of $\lambda$ over $\mathbb{Q}$, and we assume it to be of degree $n$. But $q(t)\in\mathbb{Z}[t]$ is a nonzero polynomial of degree at most n1n-1 with root $\lambda$, which is impossible by the definition of the minimal polynomial. Hence, $\chi_{M}$ and $\chi_{N}$ have no common root $\lambda\in\mathbb{C}$. ∎
@@
be square integer matrices of size $n$ and $m$, respectively, with $a_{1},b_{1}\geq 1$. Furthermore, let $\alpha,\beta\neq 0$ such that
@@
is a matrix with integer coefficients. If $M$ is nonnegative and irreducible, and if $\chi_{A},\chi_{B}\in\mathbb{Z}[t]$ are irreducible and distinct, then the characteristic polynomial of $\widetilde{M}=M+{c}y^{2}E_{11}$ is irreducible in $\mathbb{Z}[t,y]$ for all $0\neq{c}\in\mathbb{Z}$.
@@
The proof is similar to the proof of 10: the only thing to verify is that no eigenvalue of $A$ or of $B$ is also an eigenvalue of $M$. Again, we compute
@@
where $q_{1}(t)\in\mathbb{Z}[t]$ is of degree at most n1n-1 and $q_{2}(t)\in\mathbb{Z}[t]$ is of degree at most m1m-1. This is seen by developing the first column of the matrix tIMtI-M. The first coefficient is responsible for the summand $t\chi_{A}(t)\chi_{B}(t)$, the next $n$ coefficients are responsible for the summand $\pm q_{1}(t)\chi_{B}(t)$ and the final $m$ coefficients are responsible for the summand $\pm q_{2}(t)\chi_{A}(t)$. We claim that neither among $q_{1}(t)$ and $q_{2}(t)$ is the zero polynomial. Indeed, by developing the first column of the matrix $t\mathrm{I}-M$, and evaluating at $t=0$, we get
@@
which is not zero since $\chi_{A}$ is irreducible. Similarly, $q_{2}(0)\neq 0$. Now if there existed a common root $\lambda\in\mathbb{C}$ of $\chi_{M}$ and $\chi_{A}$, then $\lambda$ would also be a root of either $\chi_{B}$ or $q_{1}$. Since $\chi_{A}$ and $\chi_{B}$ are irreducible and distinct, we must have $q_{1}(\lambda)=0$. But since $\chi_{A}$ is irreducible of degree $n$, and $q_{1}(t)$ is a nonzero polynomial of degree at most n1n-1, this is impossible. Similarly, no root of $\chi_{B}$ can be a root of $\chi_{M}$, which concludes the proof. ∎
@@
One could formulate 11 with $k\geq 2$ blocks $A_{1},\dots,A_{k}$ of respective sizes $n_{1},\dots,n_{k}$, instead of $k=2$. In this case, all the $k$ characteristic polynomials $\chi_{A_{i}}$ should be irreducible and pairwise distinct. The argument is identical by considering χ_ M (t) = t ∏_i=1^k χ_A_i + ∑_i=1^k ±q_i(t) ∏_j≠i χ_A_j, where $q_{i}(t)\in\mathbb{Z}[t]$ is of degree at most $n_{i}-1$ and nonzero.
@@ 4. Warm-up: proof of Thurston’s claim
As a first illustration of our method, it is the goal of this section to provide a pair of filling multicurves $\alpha$ and $\beta$ on $S_{g}$ with multicurve intersection degree 3g33g-3. By 4, this validates Thurston’s claim that the product of two multitwists can realise the maximal possible algebraic degrees of stretch factors: 6g66g-6.
@@
We start by realising, on the surface of genus $g\geq 1$ with $2$ boundary components, a pair of filling multicurves $\alpha$ and $\beta$ such that $\chi_{XX^{\top}}\in\mathbb{Z}[t]$ is irreducible and of degree 3g13g-1. On a surface with boundary, we say that a pair of multicurves is filling if they intersect transversally and if their complement consists of boundary-parallel annuli and of discs none of which is a bigon. We proceed by induction on $g$.
@@
We consider the filling pair of multicurves $\alpha$ and $\beta$ shown in Figure 1, where one of the components of $\beta$ has y1y-1 parallel copies. Here, we think of $y$ as a variable that we specify to a value in $\mathbb{Z}_{\geq 2}$ later on.
y1y-1 copies
@@
Figure 1. Two multicurves $\alpha$ (in red) and $\beta$ (in blue) on the surface of genus one with two boundary components. The multicurve $\beta$ contains y1y-1 parallel copies of one of its components.
Observe that $X$ is a matrix of size $2\times y$ (the multicurve $\beta$ has $y$ components). We directly calculate
@@
We have $\chi_{XX^{\top}}(t)=t^{2}-(4+y)t+4(y-1)$ with discriminant $y^{2}-8y+32$, which is not a square if $y\geq 12$. Indeed, in this case we have
@@
For the inductive step, assume we have constructed on the surface of genus $g\geq 1$ with $2$ boundary components a filling pair of multicurves $\alpha^{\prime},\beta^{\prime}$ such that the characteristic polynomial $\chi^{\prime}=\chi_{XX^{\top}}\in\mathbb{Z}[t]$ is irreducible and of degree 3g13g-1. Furthermore, we inductively assume that $\alpha^{\prime}_{1}$ is a simple closed curve that encircles all the handles of the surface, as illustrated in Figure 2. Take a surface of genus 11 and two boundary components, as in the case of genus $g=1$, see Figure 1, and denote its filling pair of multicurves by $\alpha^{\prime\prime}$ and $\beta^{\prime\prime}$. We assume that the component of $\beta^{\prime\prime}$ with multiple parallel copies has b1b-1 of them. Now glue its right boundary component to the left boundary component of the surface of genus $g$, and add two new curves $\alpha_{0}$ and $\beta_{0}$ to the multicurves. The curve $\alpha_{0}$ encircles all the handles of the newly formed surface, and the curve $\beta_{0}$ twice intersects $\alpha_{0}$ but no other multicurve component. Again, see Figure 2 for an illustration. By construction, the pair of multicurves $\alpha,\beta$ is again filling.
@@
Figure 2. Two surfaces of genus 11 and $g$, respectively, and two boundary components, glued together along one of their boundary components. The curves $\alpha^{\prime}_{1}$ and $\alpha^{\prime\prime}_{1}$ are shown, each encircling all the handles of their respective surface. The new curve $\alpha_{0}$ encircles all the handles of the newly formed surface, and the new curve $\beta_{0}$ runs along the glued boundary component.
@@
$\alpha$ $=\alpha_{0}\cup\alpha^{\prime}\cup\alpha^{\prime\prime},$
$\beta$ $=\beta_{0}\cup\beta^{\prime}\cup\beta^{\prime\prime}.$
@@
Let us choose $b$ such that $\chi_{B}$ is irreducible and distinct from $\chi_{A}$. We note $a_{1}=4a$ for a positive integer $a$. Indeed, since $\alpha^{\prime}_{1}$ is separating, its geometric intersection number with each component $\beta^{\prime}_{k}$ is a multiple of two, say $2x_{k}$. In particular, we have
$a_{1}=\sum_{k}|\alpha^{\prime}_{1}\cap\beta^{\prime}_{k}|^{2}=4\sum_{k}x_{k}=4a.$
In the multicurve $\beta$, we take $y^{2}-a-1\geq 1$ parallel copies of $\beta_{0}$, for $y>0$ large enough. The matrix $XX^{\top}$ for the multicurves $\alpha$ and $\beta$ takes the form
$XX^{\top}=\left(\begin{array}[]{c|ccc|cc}4y^{2}&a_{1}&\dots&a_{n}&4&2\\ \hline\cr a_{1}&&&&&\\ \vdots&&A&&&\\ a_{n}&&&&&\\ \hline\cr 4&&&&4&2\\ 2&&&&2&b\end{array}\right).$
It is a crucial feature of the Thurston–Veech construction that for a filling pair of multicurves, such a matrix $XX^{\top}$ is always primitive and hence irreducible [Thu88, Vee89]. By 11, the polynomial $\chi_{XX^{\top}}\in\mathbb{Z}[t,y]$ is irreducible (recall that $\chi_{A}$ and $\chi_{B}$ are irreducible). Hence, by Hilbert’s irreducibility theorem, there exist infinitely many specifications of $y$ (and in particular infinitely many specifications of $y$ such that $y^{2}-a-1>0$) with $\chi_{XX^{\top}}\in\mathbb{Z}[t]$ irreducible. The degree of this polynomial equals the size of the matrix $XX^{\top}$, which equals
3g1+3=3(g+1)1,3g-1+3=3(g+1)-1,
which is exactly what we wanted to show. Finally, to justify our inductive assumption on the curve $\alpha^{\prime}_{1}$, note that $\alpha_{0}$ is again a curve that encircles all handles of the newly built surface of genus $g+1$ with two boundary components.
@@
Take any example of a filling pair of multicurves $\alpha^{\prime}$ and $\beta^{\prime}$ we constructed on the surface of genus $g-1\geq 1$ with two boundary components. In particular, the curve $\alpha^{\prime}_{1}$ encircles all the g1g-1 handles of the surface, and if
@@
is the matrix $XX^{\top}$ for the multicurves $\alpha^{\prime}$ and $\beta^{\prime}$, we have $a_{1}=4a$ and $\chi_{A}(t)$ is irreducible. We identify the two boundary components of the surface to increase the genus by one. Let $\alpha_{0}$ be a longitude of the created handle, and let $\beta_{0}$ run along the glued boundary, see Figure 3. Define the two new multicurves
@@
where we take $y^{2}-a$ copies of $\beta_{0}$.
$\cdots$
$\beta^{\prime}_{1}$
$\alpha^{\prime}_{1}$
$\beta_{0}$
$\beta_{0}$
$\alpha_{0}$
Figure 3. The closed surface of genus $g$ is obtained by identifying the left and the right boundary component of the depicted closed surface of genus g1g-1 with two boundary components. The curve $\alpha_{0}$ is a longitude of the created handle, and the curve $\beta_{0}$ runs along the glued boundary.
By construction, the pair of multicurves $\alpha,\beta$ is again filling. The matrix $XX^{\top}$ for the multicurves $\alpha$ and $\beta$ takes the form
@@
and $\chi_{XX^{\top}}\in\mathbb{Z}[t,y]$ is irreducible by 10. By Hilbert’s irreducibility theorem, there exist infinitely many specifications of $y$ (and in particular infinitely many specifications of $y$ such that $y^{2}-a>0$) with $\chi_{XX^{\top}}\in\mathbb{Z}[t]$ irreducible. The degree of this polynomial equals the size of the matrix $XX^{\top}$, which equals
3(g1)1+1=3g3.3(g-1)-1+1=3g-3.